ECH 4323: Chemical Process Control EXPERIMENTAL PROTOTYPE

Design and develop a table-top water clock. The project is completely open-ended and any working prototype will be accepted. The prototype must be built from commonly available parts and should not cost more than $30. You are not allowed to buy a kit from a hobby shop.

This is a team assignment and will be graded as follows:

Working model  3 points
Engineering and Cost Analysis 2 points
A feature that ”wows” me  2 bonus points